OpenStudy (anonymous):

Trevor's dad decided to get him a cell phone for his birthday. Cell Plus has a plan that costs $24.00 per month plus an additional $0.06 per minute. Cell Best has a plan that costs $30.00 per month plus an additional $0.05 per minute. How many minutes can Trevor talk and both cell plans cost the same amount? 600 60 300 120

OpenStudy (anonymous):

24 + (0.06x) = 30 + (0.05x)

OpenStudy (anonymous):

then what?

OpenStudy (anonymous):

Solve for x..?

OpenStudy (anonymous):

Is it 60?

OpenStudy (anonymous):

24 + (0.06x) = 30 + (0.05x) 24 + (0.06(60)) = 30 + (0.05(60)) 24 + 3.6 = 30 + 3 No.

OpenStudy (anonymous):

SOrry wrong proublem

OpenStudy (anonymous):

Simple algebra problem, see below for the work. 24 +0.06*x = 30 + .05x Subtract 24 from both sides and subtract 0.05x from both sides 0.01x = 6 Divide 6 by 0.01 x = 600

OpenStudy (anonymous):

Thank you!